1500cm3

of an unknown gas have a pressure of 10000 N/m2

at 30C° find out number of moles of gas present inside 

the container.


1
Expert's answer
2021-05-27T07:36:29-0400

P = 10000 N/m2 = 10000 Pa

V = 1500 cm3 = 0.0015 m3

n = unknown

R = 8.314 m3 Pa K−1 mol−1

T = 30°C + 273 = 303 K


Solution:

The Ideal Gas Law can be used.

The Ideal Gas Law can be expressed as: PV = nRT

To find the number of moles, solve the equation for n:

n = PV / RT

n = (10000 Pa × 0.0015 m3) / (8.314 m3 Pa K−1 mol−1 × 303 K) = 0.00595 mol = 0.006 mol

n = 0.006 mol


Answer: 0.006 mol of gas present inside the container.